OpenStudy (anonymous):

The slope of a line is 16 and the point (2, 3) is on the line. Express the line in slope-intercept form.

OpenStudy (anonymous):

16(x,y)(2,3) y2-y1 ----- x2-x1 y-3 = 16 ---- x-2 y-3=16(x-2) y-3=16x-32 y=16x-32+3 y=16x-29